Constants Part A A model of a helicopter rotor has four blades, each of length 3.80 m from the central shaft to the blade tip. The model is rotated in a wind tunnel at a rotational speed of 460 rev/min What is the linear speed of the blade tip? ΑΣ 58.26 m/s Submit Previous Answers Request Answer X Incorrect; Try Again; 6 attempts remaining Part B What is the radial acceleration of the blade tip expressed as a multiple of the acceleration of gravity, g? ΑΣ 9 Submit Request AnswerConstants Part A City A lies directly west of city B. When there is no wind, an airliner makes the round trip flight of distance s between them in a flight time of ti while traveling at the same speed in both directions When a strong, steady wind is blowing from west to east and the airliner has the same air speed as before, the round trip takes a flight time of t2 How fast is the wind blowing? km/h Submit Previous Answers Request Answer Incorrect. Try Again: 6 attempts remaining The correct answer does not depend on: t
Answer #1

Solution)

Part a) Angular velocity, w= 460*(2*3.14/60)= 48.14 rad/sec

Length of blade, r=3.8 m

So, linear speed, v=rw = 3.8*48.14= 182.95 =183 m/s

======

Part b) Radial acceleration, a=v^2/r = 182.95^2/3.8 = 8808.78=8.8*10^3 m/s (Ans)

=====

Good luck!:)